# Disentangled Clothed Avatar Generation from Text Descriptions

### v1

@@ Abstract
In this paper, we introduced a novel text-to-avatar generation method that separately generates the human body and the clothes and allows high-quality animation on the generated avatar. While recent advancements in text-to-avatar generation have yielded diverse human avatars from text prompts, these methods typically combine all elements—clothes, hair, and body—into a single 3D representation. Such an entangled approach poses challenges for downstream tasks like editing or animation. To overcome these limitations, we propose a novel disentangled 3D avatar representation named Sequentially Offset-SMPL (SO-SMPL), building upon the SMPL model. SO-SMPL represents the human body and clothes with two separate meshes, but associates them with offsets to ensure the physical alignment between the body and the clothes. Then, we design an Score Distillation Sampling(SDS)-based distillation framework to generate the proposed SO-SMPL representation from text prompts. In comparison with existing text-to-avatar methods, our approach not only achieves higher exture and geometry quality and better semantic alignment with text prompts, but also significantly improves the visual quality of character animation, virtual try-on, and avatar editing. Our project page is at https://shanemankiw.github.io/SO-SMPL/
@@ 1 Introduction
Human avatars play a significant role in visual storytelling, conveying narratives through actions, expressions, and interactions with other elements in the animated world. This helps in creating a more dynamic and engaging storytelling experience for the audience. Avatars can be designed to represent diverse characters, including different genders, races, ethnicities, and abilities, making the animation more inclusive and accessible to a wider audience. This can also promote positive representation and challenge stereotypes in the media. However, the creation of high-quality avatars is typically a costly and labor-intensive endeavor, demanding the expertise of skilled 3D modelers.
In order to achieve a photorealistic, animatable, and customizable avatar, it is imperative to employ a disentangled representation consisting of distinct components, such as the body and clothing elements. Utilizing a disentangled human mesh facilitates enhanced customization and reusability of the avatar, enabling seamless alterations to clothing while maintaining the same underlying human body structure. Furthermore, the separate components contribute to heightened realism in animation, as different avatar parts exhibit distinct motion characteristics. For instance, human body movements can be accurately animated using forward kinematics, while cloth motion may be generated through physical simulations.
Despite the evident benefits, the majority of existing research primarily concentrates on reconstructing a singular avatar mesh that encompasses all components, derived from videos or images, consequently neglecting the potential for a disentangled representation. Only a limited number of studies, such as DELTA [26] and SCARF [25], have investigated the reconstruction of a disentangled avatar representation from video sources. However, these approaches aim to reconstruct a disentangled mesh but not facilitate the arbitrary creation of a disentangled avatar representation. As a result, the development of a comprehensive method for creating disentangled human avatars remains an unresolved challenge within the field.
Extending the existing text-to-avatar pipelines to generate disentangled human avatars is not trivial. The primary obstacle lies in not only independently generating the body and clothing components but also seamlessly combining them in a physically accurate manner while maintaining visually appealing results. Directly generating the body and clothing separately often results in misaligned components, detracting from the overall avatar quality. To overcome this issue, we introduce a novel representation called Sequentially-Offset-SMPL (SO-SMPL). SO-SMPL is based on the SMPL-X [53] as it effectively characterizes coarse human shapes within its parameter space. Then, SO-SMPL adds additional trainable offsets and appearance features to the SMPL-X model to represent the unclothed human body, similar to TADA [47]. Subsequently, to represent the clothing on the body, we build upon the human body to learn additional offsets, and appearances. This sequential representation ensures that the clothing is perfectly aligned with the human body while maintaining separate mesh encodings for each component.
@@
Leveraging the power of stable diffusion with SDS losses [62], we introduce a novel approach to generate our SO-SMPL representation from text prompts. Our method achieves disentangled avatar generation by first creating the unclothed human body and subsequently distilling the clothing components based on the underlying body structure. Experiments show that our method produces high-quality human avatars with better geometry details and appearances than baseline methods [47, 32]. The disentangled nature of our SO-SMPL representation unlocks strong customization capabilities, allowing for the seamless integration of various clothing options on the same human body. Furthermore, the avatars generated by our state-of-the-art method exhibit exceptionally realistic animation results, achieved by simulating distinct motion characteristics for both clothing and human body components.
@@ 2 Related work
Distilling diffusion for 3D tasks. Many recent works have utilized diffusion models to generate 3D objects or scenes. Some works [80, 6, 40, 57, 63, 42, 51] trained 3D diffusion models to directly generate 3D representations such as point clouds, meshes, radiance fields[56], SDFs [79], DMTets[70] or tri-planes [11]. Others designed diffusion models to generate 3D-aware 2D images or textures[72, 50, 52, 49, 78, 66, 8, 28, 55, 12]. Both lines of works rely on 3D data[20, 19, 24] or multi-view images[65, 87, 91] to train.
@@
3D Human Avatar generation. Previous works on 3D avatar generation [14, 58, 27, 96, 31, 59, 7, 95] could generate diverse human geometries or textured avatars, but the generation process could not be controlled by text descriptions. With the help of large language models like CLIP [64] and text-to-image diffusion models [30, 67], recent works [9, 35, 99, 88, 43] were able to generate high-fidelity avatars with prompts or with both image and prompt [84, 38]. Other works have attempted to generate human avatars by painting the given human meshes with generated textures [73, 90]. However, the generated avatars could not be animated and utilized in CG softwares.  [47, 32, 44, 4, 92, 37] could produce animatable avatars, but they ignored garment-human interactions, leading to unsatisfactory animation quality.
Disentangled representations. Reconstructing human body and clothes as separate geometries has long been studied in the field of computer vision [18, 39, 25, 26], but generating distinct body and clothes representations has barely been explored. Previous works have made attempts to generate 3D avatars as a combination of separate components [31] or layers [86], but they did not explicitly disentangle human body and accessories, making an individual component or layer not physically meaningful. A more recent work [33] trained a diffusion-based human generation model and distinctly-represented human body and different types of clothes as separate layers, but it did not have semantic controllability and suffers from over-smoothing textures. A concurrent work TECA [93] explored disentangling hair and head ornament from text and images while our work focuses on decomposing the avatar into a human body and clothes. Another recent work AvatarFusion [34] also separates clothes from human bodies. However, AvatarFusion adopts a post-processing strategy based on the outputs of AvatarClip [32], leading to low-quality results. Our method is natively designed for disentangled generation, which produces much better quality.
Figure 2: Illustration of our SO-SMPL representation. Two vertex-wise offsets, namely the human offset $\mathbf{O}_{\text{h}}$ and the clothes offset $\mathbf{O}_{\text{c}}$ are sequentially added in order to the SMPL-X body mesh $\mathbf{T}(\beta)$ to obtain the human body mesh $\mathbf{T}_{\text{h}}(\beta,\mathbf{O}_{\text{h}})$ and clothed human mesh $\mathbf{T}_{\text{c+h}}(\beta,\mathbf{O}_{\text{h}},\mathbf{O}_{\text{c}},\mathbf{M}_{\text{c}})$, where a vertex mask $\mathbf{M}_{\text{c}}$ is calculated to determine the clothing region. Finally, we mask the clothed human mesh with $\mathbf{M}_{\text{c}}$ and obtained the separated clothes mesh $\mathbf{T}_{\text{c}}(\beta,\mathbf{O}_{\text{h}},\mathbf{O}_{\text{c}},\mathbf{M}_{\text{c}})$.
@@ 3 Preliminaries
where $t$ is the time step in diffusion model, $x=g(\theta)$ is the rendered image using a differentiable renderer $g$, $x_{t}=x+\epsilon$ is the noise version of $x$, $\epsilon$ is a sampled gaussian noise, and $\hat{\epsilon}_{\phi}(x_{t};\mathbf{y},t)$ is the predicted noise of the diffusion model.
@@ 4 Methodology
Figure 3: An overview of our generation pipeline. Our pipeline has two stages. In Stage I, we generate a base human body model by optimizing its shape parameter $\beta$, vertex offset $\mathbf{O}_{\text{h}}$ and albedo texture $\Gamma_{\text{h}}$. A ControlNet is utilized to compute a Score Distillation Sampling(SDS) Loss conditioned on an ”A”-pose map $\mathcal{P}$ for the rendered RGB image $\mathbf{I}_{\text{h}}$ and normal map $\mathbf{N}_{\text{h}}$. In Stage II, we freeze the human body model and optimize the clothes parameters $\mathbf{M}_{c}$ and $\mathbf{O}_{\text{c}}$ along with the albedo texture MLP $\Gamma_{\text{c}}$. The rendered RGB images and normal maps of both the clothed human and the clothes are used in computing the SDS losses. In both stages, we utilized a simple Phong shading model to render images from our SO-SMPL representations.
@@ 4.1 SO-SMPL representation
Human Body Geometry. We represent the human body mesh with a densified SMPL-X parametric model introduced in Eq. 2 and we add learnable vertex-wise offsets $\mathbf{O}_{\text{h}}$ to represent details on the body, similar to recent works [47]. We fix the pose of the human body mesh to an “A” pose $\theta$ and use a default expression parameter $\psi$ during the text-to-avatar generation process so that $\theta$ and $\psi$ are omitted for simplicity. Specifically, the human body geometry $\mathbf{T}_{\text{h}}$ is
@@
where $\mathbf{O}_{\text{h}}$ is trainable vertex-wise offsets and $\beta$ is the trainable human shape parameter.
@@
where $\odot$ denotes the Hadamard product between the vertex-wise offsets and masks. We could also get the clothes geometry $\mathbf{T}_{\text{c}}$ only by masking on the clothed human mesh
@@
To enable the geneartion of diverse clothes types, we use the body-part segments of SMPL-X and pre-define vertex-level mask templates of 6 different garment types, namely long shirt, short shirt, long pants, short pants, vest, and overalls. These templates are used to initialize clothes mask $\mathbf{M}_{\text{c}}$.
@@
4.2 Rendering SO-SMPL
On both the human body mesh and the clothes mesh, we model their appearances with a simple Phong shading model where the meshes are characterized by their albedo and the lighting is simply a combination of a point light and an ambient light.
Albedo We use an MLP $\Gamma_{\text{i}}$ to represent the albedo of the mesh. The albedo $\mathbf{\rho}$ of a 3D point $\mathbf{x}$ on the target mesh could be computed as i.e.:
$\mathbf{\rho}$ $=\Gamma_{i}(\gamma(\mathbf{x});\xi_{i})$ (6)
@@ 4.2 Rendering SO-SMPL
Shading model. Given a randomly-sampled point light with a position of $\mathbf{v}$ and an intensity of $\mathbf{l}_{\text{d}}$, along with the ambient light intensity $\mathbf{l}_{\text{a}}$, the shading color $\mathbf{c}$ of a surface point $\mathbf{p}$ could computed by
@@
where $\vec{\mathbf{n}}_{\mathbf{p}}$ is the normal direction of the point $\mathbf{p}$, and $\vec{\mathbf{n}}_{l}=(\mathbf{v}-\mathbf{p})/\|\mathbf{v}-\mathbf{p}\|_{2}$ is the light direction towards the surface vertice.
@@
Discussion. This shading model helps us to avoid the shadows being baked into the albedo of the generated clothes and improves the consistency between geometry and textures, as shown in Fig. 4. This enhancement can be attributed to our SO-SMPL’s ability to compute normals with greater accuracy compared to previous methods. These precisely estimated normals facilitate our shading model’s capacity to explicitly render shadows, particularly those resulting from wrinkles on the clothes.
Rendering Clothes & Human Body. From any given camera viewpoint, we can render the images from our SO-SMPL representation using rasterization. We first determine the corresponding 3D point and its normal from $\mathbf{T}_{\text{h}+\text{c}}$ on every pixel. Then, the shading colors are computed using Eq. 7, resulting in the rendered image $\mathbf{I}_{\text{h}+\text{c}}$. In a similar fashion, we render separate images for the body ($\mathbf{I}{\text{h}}$) and the clothes ($\mathbf{I}{\text{c}}$), respectively. Concurrently, we also obtain the corresponding normal maps $\mathbf{N}_{\text{h}+\text{c}}$, $\mathbf{N}_{\text{h}}$, and $\mathbf{N}_{\text{c}}$. More details are included in the supplementary material.
@@ 4.3 SO-SMPL Generation
where $\mathbf{I}_{t}=\mathbf{I}+\epsilon$ is a noisy version of the input image, $\Theta_{\text{h}}=\{\beta,\mathbf{O}_{\text{h}},\xi_{\text{h}}\}$ is the trainable parameters during human body generation, $\{\beta,\mathbf{O}_{\text{h}}\}$ are geometry parameters, $\xi_{\text{h}}$ is the texture parameters, $\hat{\epsilon}_{\phi}(\mathbf{I}_{t};\mathbf{y}_{\text{h}},\mathcal{P},t)$ is the predicted noise given text embedding $\mathbf{y}_{\text{h}}$, noise step $t$ and the pose map $\mathcal{P}$. To ensure our generated human body mesh is unclothed, we added specific descriptions to the human prompt such as “wearing tight shorts”. Additionally, to further guide the body mesh and texture away from depicting clothing, we employed negative prompts like “loose clothes, accessories”.
@@
Albedo smoothness constraint. To avoid geometries (e.g. wrinkles and bumps) and lighting (shadows and highlights) being baked into the albedo of the model, we employed an albedo smoothness loss following [98, 17] to reduce these effects. With a small perturbation $\delta$ of the input position $\mathbf{x}$, we encourage the estimated albedo from texture MLP $\Gamma_{i}\in\{\Gamma_{\text{h}},\Gamma_{\text{c}}\}$ to be consistent, i.e.
@@
Geometry constraints. In both the human body and clothing generation, we utilized geometry constraints to ensure that the mesh is smooth and does not deviate too far from the original SMPL-X mesh. We employed a Laplacian smoothness term [41] $\mathcal{L}_{s}$ to minimize the norm of graph Laplacian. We also added an offset regularizing term $\mathcal{L}_{o}=\|\mathbf{O}_{i}\|_{2}$ for the body/clothes offset $\mathbf{O}_{i}\in\{\mathbf{O}_{\text{h}},\mathbf{O}_{\text{c}}\}$. Finally the geometry regularization los could be written as $\mathcal{L}_{\text{geo}}=\mathcal{L}_{s}+\mathcal{L}_{o}$.

### v2

@@ Abstract
In this paper, we introduce a novel text-to-avatar generation method that separately generates the human body and the clothes and allows high-quality animation on the generated avatar. While recent advancements in text-to-avatar generation have yielded diverse human avatars from text prompts, these methods typically combine all elements—clothes, hair, and body—into a single 3D representation. Such an entangled approach poses challenges for downstream tasks like editing or animation. To overcome these limitations, we propose a novel disentangled 3D avatar representation named Sequentially Offset-SMPL (SO-SMPL), building upon the SMPL model. SO-SMPL represents the human body and clothes with two separate meshes but associates them with offsets to ensure the physical alignment between the body and the clothes. Then, we design a Score Distillation Sampling (SDS)-based distillation framework to generate the proposed SO-SMPL representation from text prompts. Our approach not only achieves higher texture and geometry quality and better semantic alignment with text prompts, but also significantly improves the visual quality of character animation, virtual try-on, and avatar editing. Project page: this link.
Keywords:
avatar and clothes generation score distillation sampling
@@ 1 Introduction
Human avatars play a significant role in visual storytelling, conveying narratives through actions, expressions, and interactions with other elements in the animated world, thereby creating a dynamic and engaging experience for the audience. Avatars can be designed to represent diverse characters, including different genders, races, ethnicities, and abilities, promoting inclusivity and challenging stereotypes in the media. However, creating high-quality avatars is costly and labor-intensive, requiring skilled 3D modelers. To achieve photorealistic, animatable, and customizable avatars, it is imperative to employ a disentangled representation consisting of distinct components, such as body and clothing elements. This approach enhances customization and reusability, allowing seamless alterations to clothing while maintaining the same underlying body structure. Additionally, separate components contribute to heightened realism in animation, as different parts exhibit distinct motion characteristics. For example, human body movements can be animated using forward kinematics, while cloth motion can be generated through physical simulation.
Despite evident benefits, most existing research [39, 99] focuses on reconstructing a singular avatar mesh from videos or images [107, 73], neglecting the potential for disentangled representations. Only a few studies, like DELTA [26] and SCARF [27], explore disentangled avatar reconstruction from videos but do not facilitate the arbitrary creation of such representations. Thus, a comprehensive method for creating disentangled human avatars remains an unresolved challenge. Extending text-to-avatar pipelines to generate disentangled avatars involves not only independently generating body and clothing components but also ensuring their precise alignment for visually appealing results. To address this, we introduce Sequentially-Offset-SMPL (SO-SMPL), based on SMPL-X [59] for characterizing coarse human shapes. SO-SMPL adds trainable offsets and appearance features to represent the unclothed human body, inspired by TADA [51], and further refines these to learn clothing specifics. This sequential approach ensures perfect alignment and separation of body and clothing meshes.
@@
Leveraging the power of stable diffusion with SDS losses [71], we introduce a novel approach to generate our SO-SMPL representation from text prompts. Our method achieves disentangled avatar generation by first creating the unclothed human body and subsequently distilling the clothing components based on the underlying body structure. Experiments show that our method produces high-quality human avatars with better geometry details and appearances than baseline methods [51, 35]. The disentangled nature of our SO-SMPL representation unlocks strong customization capabilities, allowing for the seamless integration of various clothing options on the same human body. Furthermore, the avatars generated by our state-of-the-art method exhibit exceptionally realistic animation results, achieved by simulating distinct motion characteristics for both clothing and human body components. Moreover, by generating another layer of offsets, we can generate multiple-layered clothes, like a suit on a T-shirt, and each layer has a separate mesh.
@@ 2 Related work
Distilling diffusion for 3D tasks. Many recent works have utilized diffusion models to generate 3D objects or scenes. Some works [91, 6, 43, 63, 72, 45, 57] trained 3D diffusion models to directly generate 3D representations such as point clouds, meshes, radiance fields[62], SDFs [90], UDFs [108], DMTets[80] or tri-planes [11]. Others designed diffusion models to generate 3D-aware 2D images or textures[82, 56, 58, 54, 89, 76, 8, 29, 61, 12]. Both lines of works rely on 3D data[21, 20, 25] or multi-view images[75, 102, 106] to train.
@@
3D Human Avatar generation. Previous works on 3D avatar generation [15, 65, 28, 113, 34, 66, 7, 112] could generate diverse human geometries or textured avatars, but the generation process could not be controlled by text descriptions. With the help of large language models like CLIP [74] and text-to-image diffusion models [32, 77], recent works [9, 38, 116, 103, 46] were able to generate high-fidelity avatars with prompts or with both image and prompt [96, 39, 70]. Other works have attempted to generate human avatars by painting the given human meshes with generated textures [84, 105]. However, the generated avatars could not be animated and utilized in CG software.  [51, 35, 47, 4, 111, 41, 55, 101] could produce animatable avatars, but they ignored garment-human interactions, leading to unsatisfactory animation quality.
Disentangled representations. Reconstructing the human body and clothes as separate geometries has long been studied in the field of computer vision [19, 42, 27, 26, 68, 52], but generating distinct body and clothes representations has barely been explored. Previous works have made attempts to generate 3D avatars as a combination of separate components [34] or layers [100], but they did not explicitly disentangle human body and accessories, making an individual component or layer not physically meaningful. A more recent work [36] trained a diffusion-based human generation model and distinctly represented the human body and different types of clothes as separate layers, but it did not have semantic controllability and suffers from over-smoothing textures. Another effort [93] proposed to use an SDS-based pipeline to generate a NeRF-based clothing layer separated from the human body, but could not be utilized in physical simulations due to its implicit representation. A concurrent work TECA [109] explored disentangling hair and head ornament from text and images while our work focuses on decomposing the avatar into a human body and clothes. Another recent work AvatarFusion [37] also separates clothes from human bodies. However, AvatarFusion adopts a post-processing strategy based on the outputs of AvatarClip [35], leading to low-quality results. Our method is natively designed for disentangled generation, which produces much better quality.
@@ 3 Preliminaries
where $t$ is the time step in diffusion model, $x=g(\theta)$ is the rendered image using a differentiable renderer $g$, $x_{t}=x+\epsilon$ is the noise version of $x$, $\epsilon$ is a sampled gaussian noise, $\hat{\epsilon}_{\phi}(x_{t};\mathbf{y},t)$ is the predicted noise of the diffusion model, and $\mathbf{y}$ is the input text prompt.
@@ 4 Methodology
Figure 2: An overview of our generation pipeline. Our pipeline has two stages. In Stage I, we generate a base human body model by optimizing its shape parameter $\beta$, vertex offset $\mathbf{O}_{\text{h}}$ and albedo texture $\Gamma_{\text{h}}$. A ControlNet is utilized to compute a Score Distillation Sampling(SDS) Loss conditioned on an “A”-pose map $\mathcal{P}$ for the rendered RGB image $\mathbf{I}_{\text{h}}$ and normal map $\mathbf{N}_{\text{h}}$. In Stage II, we freeze the human body model and optimize the clothes parameters $\mathbf{O}_{\text{c}}$ along with the albedo texture MLP $\Gamma_{\text{c}}$. The rendered RGB images and normal maps of both the clothed human and the clothes are used in computing the SDS losses. In both stages, we utilized a simple Phong shading model to render images from our SO-SMPL representations.
@@ 4.1 SO-SMPL Representation
Human body geometry. As shown in Fig. 3, we represent the human body mesh with a densified SMPL-X parametric model introduced in Eq. 2 and we add learnable vertex-wise offsets $\mathbf{O}_{\text{h}}$ to represent details on the body, similar to recent works [51]. We fix the pose of the human body mesh to an “A” pose $\theta$ and use a default expression parameter $\psi$ during the text-to-avatar generation process so that $\theta$ and $\psi$ are omitted for simplicity. Specifically, the human body geometry $\mathbf{T}_{\text{h}}$ is
@@
where $\mathbf{O}_{\text{h}}$ is trainable vertex-wise offset and $\beta$ is trainable shape parameter.
@@
where $\odot$ denotes the Hadamard product between the vertex-wise offsets and masks. We can also get the clothes geometry $\mathbf{T}_{\text{c}}$ only by masking on the clothed human mesh
@@
To enable the generation of diverse clothes types, we use the body-part segments of SMPL-X and pre-define vertex-level mask templates of 6 different garment types, namely long shirts, short shirts, long pants, short pants, vests, and overalls. These templates are used to initialize clothes mask $\mathbf{M}_{\text{c}}$. We can also support skirts generation using the skirts template from BCNet [42].
@@
4.2 SO-SMPL Rendering
On both the human body mesh and the clothes mesh, we model appearances with a Phong shading model where the meshes are characterized by their albedo and the lighting is simply a combination of a point light and an ambient light.
Albedo. We use an MLP $\Gamma_{\text{i}}$ to represent the albedo of the mesh. The albedo $\mathbf{\rho}$ of a 3D point $\mathbf{x}$ on the target mesh is computed by
$\mathbf{\rho}$ $=\Gamma_{i}(\gamma(\mathbf{x});\xi_{i}),$ (6)
@@ 4.2 SO-SMPL Rendering
Shading model. Given a randomly-sampled point light with a position of $\mathbf{v}$ and an intensity of $\mathbf{l}_{\text{d}}$, along with the ambient light intensity $\mathbf{l}_{\text{a}}$, the shading color $\mathbf{c}$ of a surface point $\mathbf{p}$ is computed by
@@
where $\vec{\mathbf{n}}_{\mathbf{p}}$ is the normal direction of the point $\mathbf{p}$, $\vec{\mathbf{n}}_{l}=(\mathbf{v}-\mathbf{p})/\|\mathbf{v}-\mathbf{p}\|_{2}$ is the light direction towards the surface vertice, and $\rho_{\mathbf{p}}$ is the albeodo on this point.
Figure 3: Illustration of our SO-SMPL representation. Two vertex-wise offsets, namely the human offset $\mathbf{O}_{\text{h}}$ and the clothes offset $\mathbf{O}_{\text{c}}$ are sequentially added in order to the SMPL-X body mesh $\mathbf{T}(\beta)$ to obtain the human body mesh $\mathbf{T}_{\text{h}}(\beta,\mathbf{O}_{\text{h}})$ and clothed human mesh $\mathbf{T}_{\text{c+h}}(\beta,\mathbf{O}_{\text{h}},\mathbf{O}_{\text{c}},\mathbf{M}_{\text{c}})$, where a vertex mask $\mathbf{M}_{\text{c}}$ is calculated to determine the clothing region. Finally, we mask the clothed human mesh with $\mathbf{M}_{\text{c}}$ and obtain the separated clothes mesh $\mathbf{T}_{\text{c}}(\beta,\mathbf{O}_{\text{h}},\mathbf{O}_{\text{c}},\mathbf{M}_{\text{c}})$.
@@
Discussion. Incorporating this shading model greatly reduces the chance that the shadows are baked into the albedo of the generated clothes, which improves the consistency between geometry and textures, as shown in Fig. 4. This shading model associates the rendering colors with the generated normals so optimizing rendering colors also changes the normal directions, which reduces the chances of getting baked-in shadows or wrinkles.
Rendering clothes & human body. From any given camera viewpoint, we can render the images from our SO-SMPL representation using rasterization. We first determine the corresponding 3D point and its normal from $\mathbf{T}_{\text{h}+\text{c}}$ on every pixel. Then, the shading colors are computed using Eq. 7, resulting in the rendered image $\mathbf{I}_{\text{h}+\text{c}}$. In a similar fashion, we render separate images for the body ($\mathbf{I}_{\text{h}}$) and the clothes ($\mathbf{I}_{\text{c}}$), respectively. Concurrently, we also obtain the corresponding normal maps $\mathbf{N}_{\text{h}+\text{c}}$, $\mathbf{N}_{\text{h}}$, and $\mathbf{N}_{\text{c}}$. More details are included in the supplementary material.
@@ 4.3 SO-SMPL Generation
where $\mathbf{I}_{t}=\mathbf{I}+\epsilon$ is a noisy version of the input image, $\Theta_{\text{h}}=\{\beta,\mathbf{O}_{\text{h}},\xi_{\text{h}}\}$ is the trainable parameters during human body generation, $\{\beta,\mathbf{O}_{\text{h}}\}$ are geometry parameters, $\xi_{\text{h}}$ is the texture parameters, $\hat{\epsilon}_{\phi}(\mathbf{I}_{t};\mathbf{y}_{\text{h}},\mathcal{P},t)$ is the predicted noise given text embedding $\mathbf{y}_{\text{h}}$, noise step $t$ and the pose map $\mathcal{P}$. To ensure our generated human body mesh is unclothed, we add specific descriptions to the human prompt such as “wearing tight shorts”. Additionally, to further guide the body mesh and texture away from depicting clothing, we employ negative prompts like “loose clothes, accessories”.
@@
Albedo smoothness constraint. To avoid geometries (e.g. wrinkles and bumps) and lighting (shadows and highlights) being baked into the albedo of the model, we employ an albedo smoothness loss following [115, 17] to reduce these effects. With a small perturbation $\delta$ of the input position $\mathbf{x}$, we encourage the estimated albedo from texture MLP $\Gamma_{i}\in\{\Gamma_{\text{h}},\Gamma_{\text{c}}\}$ to be consistent, i.e.
@@
Geometry constraints. In both the human body and clothing generation, we utilize geometry constraints to ensure that the mesh is smooth and does not deviate too far from the original SMPL-X mesh. We employ a Laplacian smoothness term [44] $\mathcal{L}_{s}$ to minimize the norm of graph Laplacian. We also add an offset regularizing term $\mathcal{L}_{o}=\|\mathbf{O}_{i}\|_{2}$ for the body/clothes offset $\mathbf{O}_{i}\in\{\mathbf{O}_{\text{h}},\mathbf{O}_{\text{c}}\}$. Thus, the geometry regularization loss is $\mathcal{L}_{\text{geo}}=\mathcal{L}_{s}+\mathcal{L}_{o}$. More implementation details such as camera configurations are included in the supplementary material.